OpenStudy (anonymous):

simplify: ((x-2)^-1-(-1)^-1)/(x-1)??

OpenStudy (anonymous):

1/(x-2)

OpenStudy (anonymous):

if i understood your expression correctly

OpenStudy (anonymous):

how did you get to that? im not sure what to do with that fraction

OpenStudy (anonymous):

hmm (1/(x-2)+1)/(x-1) =((x-1)/(x-2))/(x-1) =1/(x-2)

OpenStudy (anonymous):

how did you get that x-1/x-2?

jhonyy9 (jhonyy9):

hi do you wan to learn resolving this ?

OpenStudy (anonymous):

sure!

jhonyy9 (jhonyy9):

ok let s go how do you begin it ?

jhonyy9 (jhonyy9):

do you know the property of exponentes ?

OpenStudy (anonymous):

somewhat

jhonyy9 (jhonyy9):

write me please your exercise newly here

jhonyy9 (jhonyy9):

can you write it ?

OpenStudy (anonymous):

um i can write it best i can like i did in the question? when i get to( (1/x-2) + 1)/(x-1) im not sure what to do after

jhonyy9 (jhonyy9):

so there is (1/x-2)+1)/(x-1) this is sure ?

jhonyy9 (jhonyy9):

write me please your exercise exactly like firstly here

jhonyy9 (jhonyy9):

for i can explain to you

OpenStudy (anonymous):

i think so....? wait im confused as to what you are asking? ((x-2)^-1-(-1)^-1)/(x-1)??

jhonyy9 (jhonyy9):

this is sure like ?

jhonyy9 (jhonyy9):

ok so

jhonyy9 (jhonyy9):

there is (x-2) on the exponent -1 what sign this ?

jhonyy9 (jhonyy9):

this is equal ?

OpenStudy (anonymous):

1/x-2

jhonyy9 (jhonyy9):

ok and (-1) on exponent -1 =?

OpenStudy (anonymous):

1

jhonyy9 (jhonyy9):

no no

jhonyy9 (jhonyy9):

if x on exponent -1=

jhonyy9 (jhonyy9):

comen i wan to learn you it

OpenStudy (anonymous):

ummm was it 1/-1?

jhonyy9 (jhonyy9):

yes right

jhonyy9 (jhonyy9):

ok now continue

jhonyy9 (jhonyy9):

there is -(-1) on exponent -1

OpenStudy (anonymous):

ok 1/(x-2) -1 / (x-1)

jhonyy9 (jhonyy9):

no no

jhonyy9 (jhonyy9):

you have wrote -1 on exponent -1 = 1/-1 = ?

OpenStudy (anonymous):

1/(x-2) +1 / (x-1)?

jhonyy9 (jhonyy9):

yes so when there are two fractions with denominators differents

jhonyy9 (jhonyy9):

so you need getting the denominator common

jhonyy9 (jhonyy9):

what will be in this case ?

OpenStudy (anonymous):

oh okay so i multiply them by a common den.

jhonyy9 (jhonyy9):

no firstly tell me what will be the denominator common

jhonyy9 (jhonyy9):

comen courage !!!

jhonyy9 (jhonyy9):

write me what you have wrote there in your last sentence

jhonyy9 (jhonyy9):

are you there again ?

jhonyy9 (jhonyy9):

dont give up now

OpenStudy (anonymous):

ahhh sorry X( umm..(x+2)...i have to go soon!

jhonyy9 (jhonyy9):

(x-2)(x-1) is the denominator common so than what you need to do ?

OpenStudy (anonymous):

multiply both fractions by that

jhonyy9 (jhonyy9):

by what ?

OpenStudy (anonymous):

(x-2)(x-1). I am positive i got it now! thankyoou sooo much for helping me! I have to go

jhonyy9 (jhonyy9):

ok if you know what you need to do in continue all is right Bye